it took sarah's mother 8 hours round trip to drive to the university and bring sara back home for spring break. if the university is 112 miles from home, find her mothers average speed.
A. 29 mph
B. 28 mph
C. 55 1/2 mph
D. 14 mph

Answer by jim_thompson5910(35256)

Each one way trip takes 8/2 = 4 hours.

d = rt

112 = r*4

112/4 = r

28 = r

r = 28

So her average speed is 28 mph.
--------------------------------------------------------------------------------------------------------------

If you need more help, email me at jim_thompson5910@hotmail.com

You can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------